A car starts from rest and accelerate for 10 second till it gains a speed 90 km/h. Thereafter it maintains that speed for next 15 minutes. Thereafter it applies brakes and comes to rest with in 10 second. (i) Calculate its acceleration during first 10 s. (ii) Displacement during entire journey.


1
Expert's answer
2021-08-02T08:42:04-0400

(i)

"a=\frac{\Delta v}{\Delta t}=\frac{25\rm \frac{ m}{ s}-0\frac{m}{s}}{10\:\rm s}=2.5\:\rm m/s^2"

(ii)

"d=d_1+d_2+d_3\\
=\frac{0+25}{2}*10+25*15*60+\frac{25+0}{2}*10=22750\:\rm m"